Diffusion Morning Report 09/20/11

Ted Berg tberg at stanford.edu
Tue Sep 20 06:10:56 PDT 2011


Good Morning All,
     Looking very good this morning. No new problems.
     Left overs:TEOS trap cleaning, Epi LLA. No other issues at this time.
					
	Date: 	9/20/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*trap cleaning* 	*6.83/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*3.76/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*.18/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*12.81/30* 	** 	**
Tylan 11-LTO 	*up* 	** 	*15.7/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*yellow* 	*lla issues* 	** 	** 	**
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*B on line pressure 817 psi* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110920/d16e8bc8/attachment.html>


More information about the diffusion mailing list